A jet airliner, moving initially at 300 Km/h to the east, suddenly enters a region where the wind is blowing at 100 Km/h toward the direction 30.00 north of east. What is the new speed and direction of the aircraft relative to the ground?
1
Expert's answer
2020-10-15T10:36:56-0400


Solution

Given angle "\theta=72^0"

Speed of airliner

"V=\sqrt{v_a^2+v_w^2+2v_av_w\cos\theta}"

Where "v_a, v_w" are speed of airliner and wind with respect to ground.

Therefore new speed of airliner ith respect to ground can be expressed as

"V=\sqrt{(300)^2+(100)^2+2\times (300)\times (100)cos30^°}"

"V=389.82Km\/h"


Direction

"tan\theta=\frac{v_w}{v_a}=\frac{100}{300}=0.33"

"\theta= 18.26^°"